### v1

# DCFormer: Efficient 3D Vision-Language Modeling with Decomposed Convolution

Abstract
Vision-language models (VLMs) align visual and textual representations, enabling high-performance zero-shot classification and image-text retrieval in 2D medical imaging. However, extending VLMs to 3D medical imaging remains computationally challenging. Existing 3D VLMs rely on Vision Transformers (ViTs), which are computationally expensive due to self-attention’s quadratic complexity, or 3D convolutions, which demand excessive parameters and FLOPs as kernel size increases. We introduce DCFormer, an efficient 3D medical image encoder that factorizes 3D convolutions into three parallel 1D convolutions along depth, height, and width. This design preserves spatial information while significantly reducing computational cost. Integrated into a CLIP-based vision-language framework, DCFormer is evaluated on CT-RATE, a dataset of 50,188 paired 3D chest CT volumes and radiology reports, for zero-shot multi-abnormality detection across 18 pathologies. Compared to ViT, ConvNeXt, PoolFormer, and TransUNet, DCFormer achieves superior efficiency and accuracy, with DCFormer-Tiny reaching 62.0% accuracy and a 46.3% F1-score while using significantly fewer parameters. These results highlight DCFormer’s potential for scalable, clinically deployable 3D medical VLMs. Our codes will be publicly available.
Introduction
Deep learning has revolutionized medical imaging, enabling automated disease diagnosis, accurate prognosis, and personalized treatment planning [1, 2, 3, 4]. With the emergence of Convolutional Neural Networks (CNNs) [5, 6] and Vision Transformers (ViTs) [7], deep learning has achieved state-of-the-art performance in supervised tasks such as lesion detection and organ segmentation [8, 9, 10, 11, 12]. Despite these advancements, supervised models require training on large-scale manually annotated datasets and lack generalization across diverse medical imaging tasks and modalities.
To mitigate the reliance on labeled data, self-supervised learning (SSL) methods such as masked autoencoders [13] have emerged as powerful techniques for extracting meaningful representations from unlabeled images [14, 15]. Despite their advantages, these methods still require supervised fine-tuning on small labeled datasets. Vision-language models (VLMs) such as CLIP (Contrastive Language-Image Pretraining) [16] offer a promising alternative by aligning visual and textual representations in a shared latent space, enabling zero-shot capabilities. Initially successful in natural image domains [17, 18], CLIP has since been adapted for various 2D medical imaging tasks, including zero-shot classification, prompt-driven segmentation, image-text retrieval, radiology report generation, and visual question answering [19, 20, 21, 22, 23, 24, 25].
While CLIP-based approaches have advanced 2D medical imaging, their application to 3D imaging remains underexplored due to several key challenges. A major challenge is the scarcity of large-scale, open-source datasets containing 3D image volumes paired with textual reports. To address this, Hamamci et al. [26] introduced CT-RATE, a dataset consisting of 50,188 reconstructed 3D chest computed tomography (CT) volumes from 25,692 scans of 21,304 patients, along with their corresponding radiology reports. They also proposed CT-CLIP, a 3D VLM for chest CT that employs CT-ViT [27], a model incorporating a two-stage spatial and temporal transformer. While CT-CLIP performs well in zero-shot detection and image-text retrieval tasks, its use of self-attention mechanisms leads to significant computational overhead. Additionally, its image encoder follows a ViT-like patching strategy, directly downsampling images with a fixed patch size of 20. This strategy leads to feature loss, as it fails to extract hierarchical representations—an essential aspect of medical imaging. Prior studies have shown that hierarchical feature extraction is crucial for capturing spatial and contextual information in medical images [9, 10, 28]. Despite these limitations, CT-CLIP marks a significant step toward adapting foundation models for 3D medical imaging.
Another challenge is the computational complexity of adapting CLIP-like VLMs to 3D imaging modalities such as CT and magnetic resonance imaging (MRI). The high-resolution, volumetric nature of 3D medical data imposes substantial computational demands, particularly with standard ViT-based or 3D CNN-based encoders. While ViTs effectively capture global spatial relationships, their computational complexity scales quadratically with input size, making them inefficient for 3D volumes. Likewise, 3D convolutions, which are well-suited for capturing local spatial features, become prohibitively expensive with large kernel sizes ($k>3$). To mitigate this, depthwise convolutions [29] have been explored as an alternative, reducing both parameter count and computational cost. However, they can still be inefficient for 3D imaging when large kernels and high output channel dimensions are required. Figure 1 compares the parameters and FLOPs of 2D and 3D standard versus depthwise convolutions, showing that standard 3D convolutions demand significantly more resources than their 2D counterparts. Although depthwise convolutions improve efficiency, their computational cost scales cubically with kernel size, posing challenges for large 3D medical datasets where large receptive fields are needed to capture complex anatomical structures. Additionally, their cumulative use throughout the network significantly increases the overall computational burden.
Figure 1: FLOP and parameter comparison of 2D and 3D convolutions. For simplicity, we set the number of output channels $C=32$ and depth dimension $D=32$
To overcome these challenges of existing 3D VLMs, we propose DCFormer, a family of architectures designed for 3D medical imaging. DCFormer introduces decomposed convolutions as an alternative to self-attention and conventional 3D convolutions, significantly reducing computational complexity while preserving the ability to extract both local and global spatial features. By factorizing 3D convolutions into three 1D convolutions along each spatial dimension, DCFormer aims to balance efficiency and performance, significantly reducing model parameters and FLOPs (Figure 2). We integrate DCFormer into the CLIP framework, enabling efficient joint vision-language learning for 3D medical imaging. Our experiments demonstrate that DCFormer achieves competitive zero-shot performance on CT-RATE dataset while significantly reducing computational costs compared to existing state-of-the-art methods. We anticipate that DCFormer will serve as a stepping stone for advancing scalable and efficient image encoders for vision-language understanding and inspire further research into lightweight, high-performance architectures for clinical applications.
Figure 2: FLOP and parameter comparison between 3D depthwise convolution and decomposed depthwise convolution.
Results
We train a CLIP-based joint vision-language learning framework (see Developing a CLIP framework using DCFormer in the Methods section) using the CT-RATE dataset, consisting of 50,188 non-contrast 3D chest CT volumes paired with corresponding radiology text reports. For the text encoder, we use CXR-BERT [30]. For the image encoder, we comprehensively evaluate our proposed DCFormer architecture by benchmarking its zero-shot performance against several state-of-the-art models, including ViT [7], CT-ViT [27, 26], TransUNet [9], ConvNeXt [31], InceptionNeXt [32], and PoolFormer [33] (see SOTA image encoders in the Methods section).
model variant params (M) GFLOPS accuracy (%) f1 score (%) precision (%) recall (%)
DCFormer nano 920.3K 34.21 60.4 41.9 27.2 62.8
convnext [31] nano 3.19M 31.92 62.2 39.4 26.7 55.1
poolformer [33] nano 2.79M 27.14 60.2 37.0 24.8 52.3
DCFormer naïve 5.85M 49.48 63.1 44.5 29.5 65.5
vit [7] naïve 11.10M 39.05 55 42.5 25.8 71.5
convnext [31] naive 15.63M 96.84 60.7 42.4 27.7 63.8
poolformer [33] naïve 11.31M 63.75 60.1 39.1 25.7 56.8
transunet [9] naïve 12.48M 118.9 58.6 41.4 26.5 56.0
DCFormer tiny 15.1M 168.2 62.0 46.3 29.7 70.1
vit [7] tiny 26.34M 86.43 61 43.2 28.0 64.8
convnext [31] tiny 31.59M 156.31 62.5 42.1 28.2 60.1
transunet [9] tiny 23.93M 207.5 61.5 35.8 24.7 48.7
poolformer [33] tiny 20.68M 117.46 61.8 38.3 26.0 53.5
ctvit [26] - 101.1M 160.5 62.9 44.3 29.3 65.7
Table 1: Zero-shot performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
Zero-shot multi-abnormality detection results
Once trained to maximize the similarity between image and text embeddings, the CLIP framework enables multi-abnormality detection by inputting each abnormality as a text prompt (see *Zero-shot multi-abnormality detection with DCFormer* in the Methods section). Specifically, we use the prompts "Abnormality is present." and "Abnormality is not present." for each of the 18 distinct abnormalities, following the ablation study by Hamamci et al. [26]. We then compute the normalized probability of each abnormality being present in the CT image and evaluate model performance using accuracy, F1 score, precision, and recall.
Table 1 compares the zero-shot performance and computational efficiency of multiple models on the CT-RATE dataset for image volumes of size 512×512×256. DCFormer provides a family of models with configurations ranging from nano to tiny variants, with parameter counts between 920.3K and 15.1M and GFLOPs spanning 34.21G to 168.2G. As shown in Table 1, DCFormer consistently achieves superior accuracy, F1 score, precision, and recall across all configurations while maintaining computational efficiency.
For instance, the nano variant achieves higher F1 score, precision, and recall than ConvNeXt and PoolFormer, despite having fewer parameters and similar FLOPs. Similarly, the naïve variant outperforms ConvNeXt, PoolFormer, and TransUNet, demonstrating the efficiency and effectiveness of DCFormer’s architecture. Notably, DCFormer achieves these results with substantially fewer computational resources compared to models like TransUNet and CTViT, which require much larger parameter counts and FLOPs. For example, the naive variant of DCFormer delivers 63.1% accuracy, 44.5% F1 score, 29.5% precision, and 65.5% recall using just 5.85M parameters and 49.48 GFLOPs—far fewer parameters and similar FLOPs compared to CTViT’s 101.1M parameters and 160.5 GFLOPs and TransUnet’s 23.93M parameters and 207.5 GFLOPs.
This efficiency enables DCFormer to achieve robust performance with a simple yet effective architecture, demonstrating its scalability and suitability for applications requiring both high accuracy and computational efficiency. Overall, for zero-shot implementation, DCFormer outperforms other models in efficiency, with its lightweight variants achieving competitive results while significantly reducing computational overhead. DCFormer’s ability to outperform SOTA models with fewer parameters and FLOPs highlights its potential for applications requiring both high performance and computational efficiency.
Discussion
In this paper, we introduce DCFormer, a novel vision encoder designed to efficiently process 3D medical images within a CLIP-based vision-language framework. DCFormer addresses key computational challenges in 3D CT analysis, where existing models struggle with high-resolution volumetric data due to the quadratic scaling of self-attention in vision transformers and the high computational cost of 3D convolutions. To mitigate these issues, DCFormer leverages decomposed 3D convolutions as a token mixer, reducing FLOPs and parameter count while preserving strong spatial feature extraction capabilities. This design enables superior performance in zero-shot predictions.
We evaluate DCFormer on CT-RATE, a large-scale dataset containing 50,188 reconstructed 3D chest CT volumes paired with radiology reports. The model is tested under the zero-shot setting, benchmarking its performance against state-of-the-art methods including ViT, ConvNeXt, PoolFormer, and TransUNet. Our primary objective is to develop a model with strong zero-shot capabilities for multi-abnormality detection, minimizing the need for task-specific fine-tuning. As shown in Table 1, DCFormer consistently outperforms competing models in accuracy, F1 score, precision, and recall, all while requiring significantly fewer parameters and computational resources. This efficiency is driven by decomposed convolutions, which factorize 3D convolutions into multiple 1D operations along each spatial dimension, striking an optimal balance between performance and efficiency. Unlike CT-ViT and other computationally intensive 3D models, DCFormer offers a lightweight and scalable solution for large-scale 3D medical imaging, making it particularly well-suited for clinical applications requiring real-time decision-making.
While DCFormer demonstrates promising results, several future directions could enhance its capabilities. First, advanced prompt engineering may further improve zero-shot performance by better aligning image-text embeddings for classification tasks. Second, incorporating larger and more diverse datasets across different medical imaging modalities—such as MRI and PET—could improve generalization and further validate DCFormer’s effectiveness. Finally, integrating DCFormer with other VLMs, such as LLaVA [34] and LISA [35], could expand its applications to visual question answering and prompt-driven 3D image segmentation.
Methods
Processing 3D CT images presents significant computational challenges due to their high resolution and volumetric complexity. 3D convolution is computationally expensive due to its cubic scaling with kernel size, and the self-attention mechanism further amplifies the burden by scaling quadratically with input size. To address these challenges, we propose DCFormer, a hybrid architecture that decomposes 3D convolution into multiple 1D components, significantly reducing computational overhead. DCFormer adopts MetaFormer [33, 36] as its backbone and employs decomposed convolutions as the token mixer for feature extraction. Its hierarchical structure further enhances feature representation in 3D CT images. Finally, the DCFormer-based image encoder is integrated into a CLIP framework for image-text alignment. As demonstrated in our results, this decomposition strategy substantially reduces parameter count and computational cost while maintaining high performance in both zero-shot and fine-tuning settings.
Formulation of DCFormer block
The core concept, MetaFormer [33, 36], is a general architecture abstracted from transformers [37]. In the Metaformer architecture, the token mixer (e.g. self-attention, depthwise convolution) is not specified, while normalization, a channel mixer (e.g. MLP), and residual connections [38] are retained. The input $I$ is first passed into a patch embedding block, such as a convolution:
$X=\text{PatchEmbed}(I)=Flatten(\text{Conv}^{(k,s,p)}_{C_{0}\rightarrow C}(I))$ (1)
where $X\in\mathbb{R}^{N\times C}$ denotes the embedded tokens with sequence length $N$ and embedding dimension $C$. Here, $C_{0},k,s$ and $p$ represent input image channels, kernel size, stride and padding, respectively. These embedded tokens are then fed into the Metaformer architecture:
$X^{\prime}$ $=X+\text{TokenMixer}(\text{Norm}_{1}(X)),$ (2)
$X^{\prime\prime}$ $=X^{\prime}+\sigma(\text{Norm}_{2}(X^{\prime})W_{1})W_{2}.$ (3)
Here, $\text{Norm}_{1}$ and $\text{Norm}_{2}$ are typically batch normalization [39] or layer normalization [40]. The TokenMixer serves as the core module for spatial information interaction, $W_{1}$ and $W_{2}$ are learnable weights in a two-layer channel MLP, and $\sigma$ is a non-linear activation [41, 42].
To further enhance the MetaFormer architecture, we introduce Decomposed Convolution as the token mixer within the DCFormer block. This design leverages the computational efficiency of decomposed 1D convolutional operations along each spatial axis (height, width, and depth). By splitting the 3D convolution into three parallel 1D convolutions, Decomposed Convolution captures spatial features while significantly reducing the number of parameters and computational cost. Thus, the DCFormer block integrates decomposed convolutions as a lightweight yet powerful token mixer. Let $X\in\mathbb{R}^{B\times C\times H\times W\times D}$ denote the input feature map, where $B$ is the batch size, $C$ is the number of channels, and $H$, $W$, and $D$ represent the spatial dimensions (height, width, and depth), respectively. The decomposed convolution consists of three 1D depthwise convolutions, processing the input tensor along each spatial axis:
$X^{*}_{\text{h}}$ $=\text{DWConv}^{k_{h}\times 1\times 1}_{C\rightarrow C}(X),$ (4)
$X^{*}_{\text{w}}$ $=\text{DWConv}^{1\times k_{w}\times 1}_{C\rightarrow C}(X),$ (5)
$X^{*}_{\text{d}}$ $=\text{DWConv}^{1\times 1\times k_{d}}_{C\rightarrow C}(X)$ (6)
where ($k_{h},k_{w},k_{d}$) represents the kernel sizes in height, width and depth dimensions, respectively. In our implementation, we set $k_{h}=k_{w}=k_{d}=k\in\{13,11,9,7\}$ to leverage large kernels while maintaining computational efficiency through decomposition. After applying decomposed convolutions along each spatial axis, we normalize the resulting features separately and then combine them using elementwise summation to form the DCFormer block:
$X^{\prime}$ $=X+\text{Norm}_{h}(X^{*}_{\text{h}})+\text{Norm}_{w}(X^{*}_{\text{w}})+\text{Norm}_{d}(X^{*}_{\text{d}})$ (7)
An illustration of the DCFormer and its Pytorch-like implementation are shown in Figure 3 and Algorithm 1, respectively.
Figure 3: Block illustration of MetaNeXt, ConvNeXt and DCFormer.
class DecompConv3D(nn.Module):
def __init__(self, dim, kernel_size, stride=1, norm=True, act=None):
super().__init__()
# Initialize final activation layer
self.act = act
# Initialize normalization layer
self.norm = nn.ModuleList([nn.BatchNorm3d(dim) if norm else nn.Identity() for _ in range(3)])
# Initialize decomposed convolution layers
self.c1 = nn.Sequential(
nn.Conv3d(dim, dim, kernel_size=(kernel_size, 1, 1),
padding=(kernel_size//2, 0, 0), stride=stride, groups=dim),
self.norm[0])
self.c2 = nn.Sequential(
nn.Conv3d(dim, dim, kernel_size=(1, kernel_size, 1),
padding=(0, kernel_size//2, 0), stride=stride, groups=dim),
self.norm[1])
self.c3 = nn.Sequential(
nn.Conv3d(dim, dim, kernel_size=(1, 1, kernel_size),
padding=(0, 0, kernel_size//2), stride=stride, groups=dim),
self.norm[2])
def forward(self, x):
x = self.c1(x) + self.c2(x) + self.c3(x)
if self.act is not None:
x = self.act(x)
return x
Algorithm 1 PyTorch-style pseudocode for DecompConv3D
Complexity analysis
Given an input tensor $X\in\mathbb{R}^{1\times C\times H\times W\times D}$ with batch size $B=1$ a 3D depthwise convolution [29] with a kernel size ($k_{h},k_{w},k_{d}$) where $k_{h}=k_{w}=k_{d}=k$, the total parameter count and FLOPs can be calculated as $Ck^{3}$ and $2CHWDk^{3}$, respectively (bias is omitted for simplicity). Note that both parameters and FLOPS increase cubically as $k$ increases. This introduces extensive computational burden, especially for large kernel sizes and when these convolutions are used repeatedly in an end-to-end network, where the effect of the cubic scaling compounds over multiple layers.
By decomposing the 3D depthwise convolution into three 1D depthwise convolutions, the parameter count and FLOPS are significantly reduced to $3Ck$ and $6CHWDk$, respectively, as each 1D convolution operates independently along a single spatial dimension rather than all three simultaneously. A comparison of FLOPs and parameter counts for 3D depthwise and decomposed convolutions is shown in Figure 2. For the decomposed convolution, the computational cost scales linearly with respect to kernel size $k$, as opposed to the cubic scaling in the 3D depthwise convolution.
When the kernel size is larger than 3, the complexity increase becomes much more significant. Large kernel sizes such as 7, 11 [31, 32] and even up to 51 [43, 44] have been shown to improve model performance. In 3D imaging, our decomposition approach becomes crucial, particularly in deep networks where multiple convolution layers are applied.
Figure 4: Overview of DCFormer
Building a vision encoder with DCFormer
Based on the above findings, we formulate DCFormer as the token mixer and develop a family of architectures of varying sizes, using DCFormer as the main building block. Figure 4 illustrates the general architecture of our proposed DCFormer. Following previous work [38, 45, 31, 33, 46, 32], DCFormer adopts a hierarchical structure with four stages and an initial stem stage following [46, 47].
In the stem stage, we use a decomposed convolution with a kernel size of 7 and a stride of 4. The downsampled images are then processed with three additional decomposed convolutions, each with a kernel size of 3 and a stride of 1. The outputs of the stem stage are passed into the hierarchical structure, where each stage produces tokens of sizes $\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$, $\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$, $\frac{H}{32}\times\frac{W}{32}\times\frac{D}{32}$, and $\frac{H}{64}\times\frac{W}{64}\times\frac{D}{64}$, respectively, where $H$, $W$, $D$ represent height, width and depth of the input image volume.
We propose four model variants: nano, naive, tiny, small. The number of layers in each stage are [1, 1, 1, 1] for nano, [2, 2, 2, 2] for naive, [2, 3, 3, 2] for tiny and [2, 3, 6, 2] for small models. Additionally, each stage begins with patch embedding where a 3D max pooling operation with a kernel size of 3 and a stride of 2 is used to reduce the image dimensions [47]. Following [7, 31, 32], we set the MLP ratio to 4, while the kernel size for the decomposed convolution is set to 7, as in [45, 31]. Further configuration details are provided in Table 2.
Stage #Tokens Layer Specification DCFormer
Nano Naive Tiny Small
0 H/4 X W/4 X D/4 STEM Embed. Dim. 32 64
#Blocks 4
Patch size (block 1) 7 x 7 x 7, stride 4
Patch size (block 2, 3, 4) 3 x 3 x 3, stride 1
1 H/8 X W/8 X D/8 Patch Embedding Patch size 3 x 3 x 3, stride 2
Embed. Dim. 32 64 96
DCFormer Block Kernel size 7
MLP Ratio 4
#Blocks 1 2
2 H/16 X W/16 X D/16 Patch Embedding Patch size 3 x 3 x 3, stride 2
Embed. Dim. 64 128 192
DCFormer Block Kernel size 7
MLP Ratio 4
#Blocks 1 2 3
3 H/32 X W/32 X D/32 Patch Embedding Patch size 3 x 3 x 3, stride 2
Embed. Dim. 128 256 384
DCFormer Block Kernel size 7
MLP Ratio 4
#Blocks 1 2 3 6
4 H/64 X W/64 X D/64 Patch Embedding Patch size 3 x 3 x 3, stride 2
Embed. Dim. 256 512 768
DCFormer Block Kernel size 7
MLP Ratio 4
#Blocks 1 2
Parameters (M) 0.92 5.85 15.1 26.86
FLOPS(G) 34.21 49.48 168.2 202.9
Table 2: DCFormer configuration details.
Developing a CLIP framework using DCFormer
The CLIP framework aligns visual and contextual embeddings in a shared latent space through contrastive learning (Figure 5(a)). In this paper, we integrate our proposed DCFormer as the vision encoder in CLIP, serving as the primary module for efficient processing of 3D medical images. For the text encoder, we incorporate CXR-BERT [30], following the implementation in Hamamci et al.[26]. To compute similarities between images and radiology reports, we project visual and textual features into a shared 512-dimensional embedding space. Specifically, we apply global average pooling, followed by a linear projection to the output of the image encoder, while the text encoder output is transformed using a separate linear layer. Finally, visual and textual embeddings are aligned using a contrastive loss function [16]:
$\mathcal{L}_{\text{CLIP}}=-\frac{1}{B}\sum_{i=1}^{B}\left[\log\frac{\exp\left(\text{sim}(z_{i}^{t},z_{i}^{v})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(z_{i}^{t},z_{j}^{v})/\tau\right)}+\log\frac{\exp\left(\text{sim}(z_{i}^{v},z_{i}^{t})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(z_{i}^{v},z_{j}^{t})/\tau\right)}\right]$ (8)
where $\text{sim}(z_{i}^{t},z_{j}^{v})$ represents the cosine similarity between the text embedding $z_{i}^{t}$ for the $i-$th radiology report and the visual embedding $z_{j}^{v}$ for the $j-th$ 3D image, $\tau$ is the temperature parameter (set to 1), and $B$ is the number of text-image pairs in a training batch.
Unlike CT-CLIP [26], which lacks explicit multi-scale feature extraction, our DCFormer-based CLIP framework introduces a hierarchical structure that captures multi-scale features across different spatial resolutions. This design enhances representation learning and improves the alignment between image and text embeddings by preserving both global context and fine-grained details—crucial for accurate 3D CT image interpretation. Additionally, integrating DCFormer helps address the computational challenges associated with 3D CT imaging. By utilizing decomposed convolutions, our framework effectively reduces both the parameter count and computational overhead while maintaining robust feature extraction capabilities. As a result, the DCFormer-based CLIP framework is well-suited for large-scale 3D medical imaging applications.
Figure 5: CLIP a) training and b) zero-shot inference framework
Zero-shot multi-abnormality detection with DCFormer
For zero-shot multi-abnormality detection, we follow the same approach as CheXzero[21] and CT-CLIP[26]. Specifically, two similarity scores are computed using the cosine similarity between the embedding of the CT image of interest and the embeddings of a positive (’{Pathology} is present.’) prompt and a negative (’{Pathology} is not present.’) prompt. A softmax function is then applied to these similarity scores to estimate the likelihood of each abnormality being present in the given CT image (see Figure. 5 (b) for illustration). It should be noted that the zero-shot performance can potentially be improved by experimenting with different prompt variations. For instance, possible prompt templates including: ’Signs of {Pathology} are/are not detected in the image.’, ’There is/is not {Pathology}.’, ’{Pathology} is/is not observed.’ ’Indication/No indication of {Pathology}.’. Although prompt engineering plays a critical role in improving image-text alignment, we adopt the prompt structure proposed by Hamamci et al. [26] (’{ Pathology} is/is not present.’) to maintain consistency with prior work and ensure a fair comparison.
State-of-the-art image encoders
The pioneering work ViT [7] introduced the concept of processing images as sequences of patches, leveraging the self-attention mechanism for global feature extraction. However, its quadratic complexity imposes a significant computational burden, particularly for 3D medical images. TransUNet [9] was the first model to incorporate ViTs into medical image segmentation, employing a hybrid architecture that combines CNNs with transformers. ConvNeXt [31] demonstrated that 7×7 depthwise convolutions can serve as an effective token mixer, improving performance over self-attention while maintaining computational efficiency. PoolFormer [33] introduced a more efficient feature extraction mechanism than ConvNeXt, replacing self-attention with a simple average pooling operation to minimize computational overhead while preserving strong performance.
Datasets
For model training and evaluation, we used the open-source CT-RATE dataset [26], which consists of 50,188 reconstructed CT volumes from 25,692 distinct CT experiments involving 21,304 patients. Each CT volume is paired with a radiology report. The dataset also includes 18 distinct abnormalities extracted from medical reports of each CT scan (findings and impressions). Table 3 provides an in-depth overview of CT-RATE, detailing the distribution of abnormalities across training and validation subsets. Each abnormality is associated with the number of samples in both sets, along with their respective ratios within the dataset. These ratios represent the proportion of samples for each abnormality relative to the total dataset. Notably, the ratios for the training and validation sets remain nearly identical across all abnormalities, ensuring that the validation set accurately reflects the training distribution, making it reliable for evaluating model performance.
Implementation details
For model training, we follow the same data splitting and pre-processing strategy as in [26]. Specifically, we use 20,000 patients for training and 1,304 for validation. For pre-processing, we first resize each CT volume to a spacing of 0.75 mm on the x-axis and y-axis and 1.5 mm on the z-axis. Then, the volumes are center-cropped or padded to a fixed size of 512 x 512 x 256. Finally, we clip the Hounsfield Unit (HU) values of each CT image volume to [-1000,1000] and normalize them to [-1,1]. We train models using the AdamW optimizer [48] with an initial learning rate $1^{-5}$. We dot not apply learning rate scheduling or warmup, as we have not observed any significant improvements. All models are trained for 15 epochs.
Abnormality Number of samples (train) Number of samples (validation) Ratio in the dataset (train) Ratio in the dataset (validation)
Medical Material 5818 313 0.123 0.103
Arterial wall calcification 13377 867 0.284 0.285
Cardiomegaly 5308 325 0.113 0.107
Pericardial effusion 3412 226 0.072 0.074
Coronary artery wall calcification 12025 765 0.255 0.252
Hiatal hernia 6751 417 0.143 0.137
Lymphadenopathy 12221 789 0.259 0.260
Emphysema 9122 600 0.193 0.197
Atelectasis 12263 713 0.260 0.235
Lung nodule 21382 1361 0.453 0.448
Lung opacity 17420 1184 0.369 0.390
Pulmonary fibrotic sequela 12589 831 0.267 0.273
Pleural effusion 5705 376 0.121 0.124
Mosaic attenuation pattern 3638 253 0.077 0.083
Peribronchial thickening 4973 355 0.105 0.117
Consolidation 8319 581 0.176 0.191
Bronchiectasis 4732 330 0.100 0.109
Interlobular septal thickening 3745 249 0.079 0.082
Table 3: Detailed overview of the CT-RATE dataset for each abnormality and their distributions in the training and validation sets. [26]
Acknowledgements
This work was supported by the Department of Medicine and the Intelligent Clinical Care Center at the University of Florida College of Medicine.
Author contributions statement
G.C.A. conducted study design, software development, experimental setup, data analysis, visualization, and manuscript writing. K.G. provided numerous insights and contributed to discussions throughout this study. W.S. oversaw the project, conducted study design, experimental setup, and data analysis, and contributed to manuscript writing. All authors reviewed the manuscript.
Additional information
Competing interests The authors have no conflict of interest to declare.

### v2

@@
-DCFormer: Efficient 3D Vision-Language Modeling with Decomposed Convolution
+DCFormer: Efficient 3D Vision-Language Modeling with Decomposed Convolutions
 Abstract
-Vision-language models (VLMs) align visual and textual representations, enabling high-performance zero-shot classification and image-text retrieval in 2D medical imaging.
-However, extending VLMs to 3D medical imaging remains computationally challenging.
-Existing 3D VLMs rely on Vision Transformers (ViTs), which are computationally expensive due to self-attention’s quadratic complexity, or 3D convolutions, which demand excessive parameters and FLOPs as kernel size increases.
-We introduce DCFormer, an efficient 3D medical image encoder that factorizes 3D convolutions into three parallel 1D convolutions along depth, height, and width.
+Vision-language models (VLMs) have been widely applied to 2D medical image analysis due to their ability to align visual and textual representations.
+However, extending VLMs to 3D imaging remains computationally challenging.
+Existing 3D VLMs often rely on Vision Transformers (ViTs), which are computationally expensive due to the quadratic complexity of self-attention, or on 3D convolutions, which require large numbers of parameters and FLOPs as kernel size increases.
+We introduce DCFormer, an efficient 3D image encoder that factorizes 3D convolutions into three parallel 1D convolutions along the depth, height, and width dimensions.
 This design preserves spatial information while significantly reducing computational cost.
-Integrated into a CLIP-based vision-language framework, DCFormer is evaluated on CT-RATE, a dataset of 50,188 paired 3D chest CT volumes and radiology reports, for zero-shot multi-abnormality detection across 18 pathologies.
-Compared to ViT, ConvNeXt, PoolFormer, and TransUNet, DCFormer achieves superior efficiency and accuracy, with DCFormer-Tiny reaching 62.0% accuracy and a 46.3% F1-score while using significantly fewer parameters.
+Integrated into a CLIP-based vision-language framework, DCFormer is trained and evaluated on CT-RATE, a dataset of 50,188 paired 3D chest CT volumes and radiology reports.
+In zero-shot and fine-tuned detection of 18 pathologies, as well as in image–text retrieval tasks, DCFormer consistently outperforms state-of-the-art 3D vision encoders, including CT-ViT, ViT, ConvNeXt, PoolFormer, and TransUNet.
 These results highlight DCFormer’s potential for scalable, clinically deployable 3D medical VLMs.
-Our codes will be publicly available.
+Our code is available at: https://github.com/mirthAI/DCFormer.
 Introduction
-Deep learning has revolutionized medical imaging, enabling automated disease diagnosis, accurate prognosis, and personalized treatment planning [1, 2, 3, 4].
-With the emergence of Convolutional Neural Networks (CNNs) [5, 6] and Vision Transformers (ViTs) [7], deep learning has achieved state-of-the-art performance in supervised tasks such as lesion detection and organ segmentation [8, 9, 10, 11, 12].
-Despite these advancements, supervised models require training on large-scale manually annotated datasets and lack generalization across diverse medical imaging tasks and modalities.
-To mitigate the reliance on labeled data, self-supervised learning (SSL) methods such as masked autoencoders [13] have emerged as powerful techniques for extracting meaningful representations from unlabeled images [14, 15].
-Despite their advantages, these methods still require supervised fine-tuning on small labeled datasets.
+Deep learning has revolutionized medical imaging by enabling automated disease diagnosis, accurate prognosis, and personalized treatment planning [1, 2, 3, 4].
+In particular, Convolutional Neural Networks (CNNs) [5, 6] and Vision Transformers (ViTs) [7] have driven state-of-the-art performance in many supervised tasks (e.g., lesion detection and organ segmentation [8, 9, 10, 11, 12]).
+However, these supervised methods require training on large-scale manually annotated datasets and often fail to generalize across different imaging modalities and tasks.
+This limitation has motivated the development of more scalable and generalizable learning paradigms.
+To reduce the dependency on extensive data labeling, self-supervised learning (SSL) methods such as masked autoencoders [13] have been developed to extract informative representations from unlabeled images [14, 15].
+While effective, these SSL methods still typically require subsequent supervised fine-tuning on small labeled datasets.
 Vision-language models (VLMs) such as CLIP (Contrastive Language-Image Pretraining) [16] offer a promising alternative by aligning visual and textual representations in a shared latent space, enabling zero-shot capabilities.
-Initially successful in natural image domains [17, 18], CLIP has since been adapted for various 2D medical imaging tasks, including zero-shot classification, prompt-driven segmentation, image-text retrieval, radiology report generation, and visual question answering [19, 20, 21, 22, 23, 24, 25].
-While CLIP-based approaches have advanced 2D medical imaging, their application to 3D imaging remains underexplored due to several key challenges.
-A major challenge is the scarcity of large-scale, open-source datasets containing 3D image volumes paired with textual reports.
-To address this, Hamamci et al.
-[26] introduced CT-RATE, a dataset consisting of 50,188 reconstructed 3D chest computed tomography (CT) volumes from 25,692 scans of 21,304 patients, along with their corresponding radiology reports.
-They also proposed CT-CLIP, a 3D VLM for chest CT that employs CT-ViT [27], a model incorporating a two-stage spatial and temporal transformer.
-While CT-CLIP performs well in zero-shot detection and image-text retrieval tasks, its use of self-attention mechanisms leads to significant computational overhead.
-Additionally, its image encoder follows a ViT-like patching strategy, directly downsampling images with a fixed patch size of 20.
-This strategy leads to feature loss, as it fails to extract hierarchical representations—an essential aspect of medical imaging.
-Prior studies have shown that hierarchical feature extraction is crucial for capturing spatial and contextual information in medical images [9, 10, 28].
-Despite these limitations, CT-CLIP marks a significant step toward adapting foundation models for 3D medical imaging.
-Another challenge is the computational complexity of adapting CLIP-like VLMs to 3D imaging modalities such as CT and magnetic resonance imaging (MRI).
-The high-resolution, volumetric nature of 3D medical data imposes substantial computational demands, particularly with standard ViT-based or 3D CNN-based encoders.
-While ViTs effectively capture global spatial relationships, their computational complexity scales quadratically with input size, making them inefficient for 3D volumes.
-Likewise, 3D convolutions, which are well-suited for capturing local spatial features, become prohibitively expensive with large kernel sizes ($k>3$).
+Initially developed for natural image domains [17, 18], CLIP has since been adapted for 2D medical imaging tasks, including zero-shot classification, prompt-based segmentation, image-text retrieval, radiology report generation, and visual question answering [19, 20, 21, 22, 23, 24, 25].
+Despite the success of VLMs in 2D, their application to 3D imaging remains underexplored due to several challenges.
+One major challenge is the scarcity of large-scale public datasets of 3D image volumes paired with text reports.
+Recently, Hamamci et al.
+[26] addressed this by releasing CT-RATE, a dataset consisting of 50,188 reconstructed 3D chest computed tomography (CT) volumes (from 25,692 scans of 21,304 patients) with corresponding radiology reports.
+They also introduced CT-CLIP, a 3D VLM for chest CT that uses a vision transformer encoder (CT-ViT [27]) with a two-stage spatial and temporal transformer architecture.
+Although CT-CLIP achieved promising results in zero-shot detection and image-text retrieval tasks, its reliance on self-attention for the image encoder leads to significant computational overhead.
+Additionally, CT-CLIP uses a ViT-like patch tokenization strategy that downsamples each volume using a fixed patch size of 20.
+This strategy causes a loss of fine-grained information, as it fails to extract hierarchical representations — an essential element in medical imaging.
+Prior work has shown that capturing multi-scale, hierarchical features is crucial for understanding complex spatial and contextual information in medical images [9, 10, 28].
+Nonetheless, CT-CLIP marks a significant step toward bringing VLMs to 3D medical imaging.
+Figure 1: Parameter count and computational cost (FLOPs) comparison for 2D and 3D standard and depthwise convolutions across kernel sizes.
+For simplicity, the number of output channels is fixed at $C=32$, and the depth dimension is set to $D=32$.
+Another barrier to 3D VLMs is the computational complexity of processing volumetric data (e.g., CT scans).
+High-resolution 3D image volumes impose heavy computational demands on standard ViT-based or 3D CNN-based encoders.
+While ViTs are effective at capturing global relationships, their computational complexity scales quadratically with input size, making them inefficient for 3D volumes.
+Similarly, 3D convolutions are well-suited for local feature extraction, but they can become prohibitively expensive as convolutional kernel size grows beyond 3.
 To mitigate this, depthwise convolutions [29] have been explored as an alternative, reducing both parameter count and computational cost.
@@
 Figure 1 compares the parameters and FLOPs of 2D and 3D standard versus depthwise convolutions, showing that standard 3D convolutions demand significantly more resources than their 2D counterparts.
-Although depthwise convolutions improve efficiency, their computational cost scales cubically with kernel size, posing challenges for large 3D medical datasets where large receptive fields are needed to capture complex anatomical structures.
-Additionally, their cumulative use throughout the network significantly increases the overall computational burden.
-Figure 1: FLOP and parameter comparison of 2D and 3D convolutions. For simplicity, we set the number of output channels $C=32$ and depth dimension $D=32$
-To overcome these challenges of existing 3D VLMs, we propose DCFormer, a family of architectures designed for 3D medical imaging.
-DCFormer introduces decomposed convolutions as an alternative to self-attention and conventional 3D convolutions, significantly reducing computational complexity while preserving the ability to extract both local and global spatial features.
-By factorizing 3D convolutions into three 1D convolutions along each spatial dimension, DCFormer aims to balance efficiency and performance, significantly reducing model parameters and FLOPs (Figure 2).
-We integrate DCFormer into the CLIP framework, enabling efficient joint vision-language learning for 3D medical imaging.
-Our experiments demonstrate that DCFormer achieves competitive zero-shot performance on CT-RATE dataset while significantly reducing computational costs compared to existing state-of-the-art methods.
-We anticipate that DCFormer will serve as a stepping stone for advancing scalable and efficient image encoders for vision-language understanding and inspire further research into lightweight, high-performance architectures for clinical applications.
-Figure 2: FLOP and parameter comparison between 3D depthwise convolution and decomposed depthwise convolution.
+Although depthwise convolutions improve efficiency, their computational cost scales cubically with kernel size, posing challenges for 3D medical imaging where large receptive fields are needed to capture complex anatomical structures.
+Additionally, stacking many such layers in a deep network compounds computational burden.
+To address these challenges of existing 3D VLMs, we propose DCFormer, a family of 3D vision-language encoder architectures for medical images.
+DCFormer replaces both self-attention and standard 3D convolutions with a decomposed convolution strategy that significantly reduces computational complexity while preserving the ability to extract both local and global spatial features.
+In essence, each 3D convolution in DCFormer is factorized into three 1D convolutions applied along the height, width, and depth dimensions (Figure 3).
+This factorization significantly reduces the number of parameters and FLOPs required (Figure 2), achieving a much better balance of efficiency and performance.
+We integrate DCFormer as the vision encoder into a CLIP framework to enable efficient image-text representation learning on 3D imaging data.
+In our experiments on the CT-RATE dataset, DCFormer achieves competitive or better performance in zero-shot pathology detection and image-text retrieval tasks, while significantly reducing computational costs compared to existing state-of-the-art methods.
+To our knowledge, DCFormer is the first 3D vision encoder to use decomposed convolutions to improve vision-language modeling.
+We anticipate that DCFormer can serve as a stepping stone toward scalable, efficient image encoders for vision-language understanding, and hope it will inspire further research into lightweight, high-performance architectures for clinical applications.
+Figure 2: Parameter count and computational cost (FLOPs) comparison for standard 3D depthwise convolution and decomposed depthwise convolution.
 Results
-We train a CLIP-based joint vision-language learning framework (see Developing a CLIP framework using DCFormer in the Methods section) using the CT-RATE dataset, consisting of 50,188 non-contrast 3D chest CT volumes paired with corresponding radiology text reports.
-For the text encoder, we use CXR-BERT [30].
-For the image encoder, we comprehensively evaluate our proposed DCFormer architecture by benchmarking its zero-shot performance against several state-of-the-art models, including ViT [7], CT-ViT [27, 26], TransUNet [9], ConvNeXt [31], InceptionNeXt [32], and PoolFormer [33] (see SOTA image encoders in the Methods section).
+We trained a CLIP-based vision-language framework (see Methods: Developing a CLIP framework using DCFormer ) using the CT-RATE dataset, which consists of 50,188 non-contrast 3D chest CT volumes paired with corresponding free-text radiology reports.
+For the text encoder, we used a domain-specific BERT model (CXR-BERT [30]).
+For the image encoder, we evaluated different variants of our proposed DCFormer architecture (nano, naïve, tiny) against several state-of-the-art 3D vision encoders, including ViT3D [7], CT-ViT [27, 26], TransUNet [9], ConvNeXt [31], and PoolFormer [32].
+We focused on three main tasks: zero-shot abnormality detection, fine-tuned abnormality detection, and image–text retrieval.
+Zero-Shot Multi-Abnormality Detection
+Once trained to align images and text in a shared latent space, our CLIP model can perform zero-shot detection of multiple pathologies in a chest CT volume by using each pathology name as a text prompt (Figure 5b)).
+We adopt an approach similar to CheXzero[21] and CT-CLIP[26]: for each of the 18 candidate abnormalities, we provide a positive (’{Pathology} is present.’) prompt and a negative (’{Pathology} is not present.’) prompt.
+We compute the cosine similarity between between the CT image embedding and each of the text prompt embeddings.
+These similarities are used to compute the normalized probability of each abnormality being present.
+We evaluate model performance using accuracy, F1 score, precision, and recall.
+Table 1: Zero-shot performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
 Model Variant Params (M) GFLOPS Accuracy (%) F1 Score (%) Precision (%) Recall (%)
-DCFormer nano 920.3K 34.21 60.4 41.9 27.2 62.8
-convnext [31] nano 3.19M 31.92 62.2 39.4 26.7 55.1
-poolformer [33] nano 2.79M 27.14 60.2 37.0 24.8 52.3
-DCFormer naïve 5.85M 49.48 63.1 44.5 29.5 65.5
-vit [7] naïve 11.10M 39.05 55 42.5 25.8 71.5
-convnext [31] naive 15.63M 96.84 60.7 42.4 27.7 63.8
-poolformer [33] naïve 11.31M 63.75 60.1 39.1 25.7 56.8
-transunet [9] naïve 12.48M 118.9 58.6 41.4 26.5 56.0
-DCFormer tiny 15.1M 168.2 62.0 46.3 29.7 70.1
-vit [7] tiny 26.34M 86.43 61 43.2 28.0 64.8
-convnext [31] tiny 31.59M 156.31 62.5 42.1 28.2 60.1
-transunet [9] tiny 23.93M 207.5 61.5 35.8 24.7 48.7
-poolformer [33] tiny 20.68M 117.46 61.8 38.3 26.0 53.5
-ctvit [26] - 101.1M 160.5 62.9 44.3 29.3 65.7
-Table 1: Zero-shot performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
-Zero-shot multi-abnormality detection results
-Once trained to maximize the similarity between image and text embeddings, the CLIP framework enables multi-abnormality detection by inputting each abnormality as a text prompt (see *Zero-shot multi-abnormality detection with DCFormer* in the Methods section).
-Specifically, we use the prompts "Abnormality is present." and "Abnormality is not present." for each of the 18 distinct abnormalities, following the ablation study by Hamamci et al.
-[26].
-We then compute the normalized probability of each abnormality being present in the CT image and evaluate model performance using accuracy, F1 score, precision, and recall.
-Table 1 compares the zero-shot performance and computational efficiency of multiple models on the CT-RATE dataset for image volumes of size 512×512×256.
-DCFormer provides a family of models with configurations ranging from nano to tiny variants, with parameter counts between 920.3K and 15.1M and GFLOPs spanning 34.21G to 168.2G.
-As shown in Table 1, DCFormer consistently achieves superior accuracy, F1 score, precision, and recall across all configurations while maintaining computational efficiency.
-For instance, the nano variant achieves higher F1 score, precision, and recall than ConvNeXt and PoolFormer, despite having fewer parameters and similar FLOPs.
-Similarly, the naïve variant outperforms ConvNeXt, PoolFormer, and TransUNet, demonstrating the efficiency and effectiveness of DCFormer’s architecture.
-Notably, DCFormer achieves these results with substantially fewer computational resources compared to models like TransUNet and CTViT, which require much larger parameter counts and FLOPs.
-For example, the naive variant of DCFormer delivers 63.1% accuracy, 44.5% F1 score, 29.5% precision, and 65.5% recall using just 5.85M parameters and 49.48 GFLOPs—far fewer parameters and similar FLOPs compared to CTViT’s 101.1M parameters and 160.5 GFLOPs and TransUnet’s 23.93M parameters and 207.5 GFLOPs.
-This efficiency enables DCFormer to achieve robust performance with a simple yet effective architecture, demonstrating its scalability and suitability for applications requiring both high accuracy and computational efficiency.
-Overall, for zero-shot implementation, DCFormer outperforms other models in efficiency, with its lightweight variants achieving competitive results while significantly reducing computational overhead.
-DCFormer’s ability to outperform SOTA models with fewer parameters and FLOPs highlights its potential for applications requiring both high performance and computational efficiency.
+DCFormer nano 0.92 34.21 60.4 41.9 27.2 62.8
+naïve 5.85 49.48 63.1 44.5 29.5 65.5
+tiny 15.1 168.2 62.0 46.3 29.7 70.1
+ConvNeXt [31] nano 3.19 31.92 62.2 39.4 26.7 55.1
+naïve 15.63 96.84 60.7 42.4 27.7 63.8
+tiny 31.59 156.31 62.5 42.1 28.2 60.1
+PoolFormer [32] nano 2.79 27.14 60.2 37.0 24.8 52.3
+naïve 11.31 63.75 60.1 39.1 25.7 56.8
+tiny 20.68 117.46 61.8 38.3 26.0 53.5
+TransUNet [9] naïve 12.48 118.9 58.6 41.4 26.5 56.0
+tiny 23.93 207.5 61.5 35.8 24.7 48.7
+ViT [7] naïve 11.10 39.05 55.0 42.5 25.8 71.5
+tiny 26.34 86.43 61.0 43.2 28.0 64.8
+CTViT [26] - 101.1 160.5 62.9 44.3 29.3 65.7
+Table 1 compares the zero-shot performance and computational efficiency of multiple models on the CT-RATE dataset (image volume size 512×512×256).
+DCFormer provides a family of models with configurations ranging from nano to tiny variants.
+Across all configurations, the lightweight variants of DCFormer consistently achieve superior results while reducing computational overhead.
+For instance, the nano variant of DCFormer achieves higher F1 score, precision, and recall than the corresponding nano variants of ConvNeXt and PoolFormer.
+Similarly, the naïve variant of DCFormer achieves a higher F1 score than the naïve variants of ConvNeXt, PoolFormer, and TransUNet.
+Notably, DCFormer achieves these results with significantly improved computational efficiency.
+For example, the naïve variant of DCFormer delivers a 44.5% F1 score—outperforming CT-ViT as well as the tiny variants of ConvNeXt, PoolFormer, and TransUNet—while using substantially fewer parameters and FLOPs: 5.85 million parameters and 49.48 GFLOPs, compared to CT-ViT’s 101.1 million parameters and 160.5 GFLOPs.
+DCFormer’s ability to outperform state-of-the-art models with far fewer computational resources underscores its potential for applications demanding both high performance and efficiency.
+Fine-Tuned Multi-Abnormality Detection
+Table 2: Fine-tuning performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
+Model Variant Params (M) GFLOPS Accuracy (%) F1 Score (%) Precision (%) Recall (%)
+DCFormer nano 0.92 34.21 60.1 45.0 28.6 69.6
+naïve 5.85 49.48 64.2 47.2 31.0 70.2
+tiny 15.1 168.2 64.1 48.6 31.4 72.9
+ConvNeXt [31] nano 3.19 31.92 56.3 43.5 26.2 70.1
+naïve 15.63 96.84 59.0 44.2 27.2 67.6
+tiny 31.59 156.31 60.4 44.8 28.2 68.2
+PoolFormer [32] nano 2.79 27.14 52.6 41.2 24.1 67.6
+naïve 11.31 63.75 53.9 41.9 24.8 68.7
+tiny 20.68 117.46 54.2 42.2 25.0 68.8
+TransUNet [9] naïve 12.48 118.9 60.2 44.6 29.4 70.1
+tiny 23.93 207.5 50.2 39.4 22.7 66.1
+ViT [7] naïve 11.10 39.05 56.2 43.1 26.1 69.8
+tiny 26.34 86.43 59.4 44.9 28.0 70.8
+CTViT [26] - 101.1 160.5 61.8 46.4 29.4 70.1
+To further validate DCFormer  we fine-tuned each model by adding a single linear classification layer on top of the frozen image encoder, and trained on labeled data for the same 18 chest abnormalities (Figure 5c).
+We used a Binary Cross-Entropy (BCE) loss to allow independent prediction of each abnormality (i.e., treating it as a multi-label classification rather than multi-class).
+Table 2 summarizes the fine-tuning results on the test set.
+As expected, fine-tuning improves the F1 scores of all models compared to their zero-shot performance.
+Notably, DCFormer variants continue to outperform all other architectures when fine-tuned.
+The DCFormer-tiny model achieves the highest F1 score at 48.6%, surpassing the best competing model (CTViT, 46.4%).
+DCFormer-naïve also achieves a strong F1 score of 47.2%, outperforming all other fine-tuned models.
+These gains highlight DCFormer’s efficient and effective feature extraction when adapted to specific tasks.
+Image-Text Retrieval
+Finally, we evaluated DCFormer on bidirectional image-text retrieval tasks.
+We computed cosine similarity scores between each CT volume embedding and each report embedding, and evaluated retrieval performance using Recall@k.
+As shown in Table 3, DCFormer delivers strong performance in both image-to-text (IR) and text-to-image (TR) tasks.
+The DCFormer tiny variant achieves the highest Recall@50, with 11.31% for IR and 12.37% for TR.
+The naïve variant of DCFormer significantly outperforms the nano variant, highlighting the effective scaling capability of our model.
+These results demonstrate DCFormer’s ability to capture semantic relationships between visual and textual data.
+We attribute the strong retrieval results to DCFormer’s use of large-kernel decomposed convolutions, which provide broad spatial context while maintaining efficiency.
+Table 3: Image-text retrieval performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
+Models Variant IR TR
+R@1 R@10 R@50 R@1 R@10 R@50
+DCFormer nano 0.0001 0.0135 0.0617 0.0001 0.0135 0.0707
+naive 0.0026 0.0220 0.0954 0.0016 0.0214 0.0971
+tiny 0.0053 0.0296 0.1131 0.0039 0.0280 0.1237
+ConvNeXt[31] nano 0.0013 0.0102 0.0562 0.0029 0.0135 0.0599
+naive 0.0016 0.0161 0.0780 0.0023 0.0214 0.0740
+tiny 0.0026 0.0135 0.0694 0.0010 0.0171 0.0796
+PoolFormer[32] nano 0.0016 0.0095 0.0480 0.0023 0.0164 0.0592
+naive 0.0013 0.0134 0.0569 0.0026 0.0151 0.0635
+tiny 0.0007 0.0171 0.0638 0.0016 0.0115 0.0648
+TransUNet[9] naive 0.0023 0.0236 0.1000 0.0033 0.0269 0.1105
+tiny 0.0016 0.0131 0.0589 0.0026 0.0151 0.0651
+ViT[7] naive 0.0016 0.0112 0.0500 0.0007 0.0109 0.0563
+tiny 0.0010 0.0167 0.0628 0.0013 0.0148 0.0737
+CTViT[26] - 0.0046 0.0293 0.1096 0.0023 0.0260 0.0921
 Discussion
-In this paper, we introduce DCFormer, a novel vision encoder designed to efficiently process 3D medical images within a CLIP-based vision-language framework.
-DCFormer addresses key computational challenges in 3D CT analysis, where existing models struggle with high-resolution volumetric data due to the quadratic scaling of self-attention in vision transformers and the high computational cost of 3D convolutions.
-To mitigate these issues, DCFormer leverages decomposed 3D convolutions as a token mixer, reducing FLOPs and parameter count while preserving strong spatial feature extraction capabilities.
-This design enables superior performance in zero-shot predictions.
-We evaluate DCFormer on CT-RATE, a large-scale dataset containing 50,188 reconstructed 3D chest CT volumes paired with radiology reports.
-The model is tested under the zero-shot setting, benchmarking its performance against state-of-the-art methods including ViT, ConvNeXt, PoolFormer, and TransUNet.
-Our primary objective is to develop a model with strong zero-shot capabilities for multi-abnormality detection, minimizing the need for task-specific fine-tuning.
-As shown in Table 1, DCFormer consistently outperforms competing models in accuracy, F1 score, precision, and recall, all while requiring significantly fewer parameters and computational resources.
-This efficiency is driven by decomposed convolutions, which factorize 3D convolutions into multiple 1D operations along each spatial dimension, striking an optimal balance between performance and efficiency.
-Unlike CT-ViT and other computationally intensive 3D models, DCFormer offers a lightweight and scalable solution for large-scale 3D medical imaging, making it particularly well-suited for clinical applications requiring real-time decision-making.
-While DCFormer demonstrates promising results, several future directions could enhance its capabilities.
-First, advanced prompt engineering may further improve zero-shot performance by better aligning image-text embeddings for classification tasks.
-Second, incorporating larger and more diverse datasets across different medical imaging modalities—such as MRI and PET—could improve generalization and further validate DCFormer’s effectiveness.
-Finally, integrating DCFormer with other VLMs, such as LLaVA [34] and LISA [35], could expand its applications to visual question answering and prompt-driven 3D image segmentation.
+In this paper, we introduced DCFormer, a novel 3D vision encoder designed to efficiently process volumetric medical images within vision–language models.
+DCFormer addresses key computational challenges in 3D image analysis, where conventional models struggle due to the quadratic complexity of vision transformers and the computational burden of large-kernel 3D convolutions.
+By leveraging decomposed 3D convolutions as a token mixing mechanism, DCFormer drastically reduces parameter count while preserving strong spatial feature extraction capabilities.
+We evaluated DCFormer on CT-RATE, a large-scale dataset comprising 50,188 3D chest CT volumes paired with corresponding radiology reports, benchmarking its performance against several state-of-the-art 3D vision encoders, including ViT, ConvNeXt, PoolFormer, TransUNet, and CT-ViT.
+Across tasks such as multi-abnormality detection (in both zero-shot and fine-tuned settings) and image–text retrieval, DCFormer consistently outperformed all competing models.
+This performance is directly attributable to its decomposed convolution architecture, which captures multi-scale spatial features without the computational burden of self-attention or full 3D convolutions.
+Unlike CT-ViT and other resource-intensive models, DCFormer achieves a favorable balance between accuracy and efficiency, making it a practical and scalable solution for volumetric medical imaging.
+Importantly, DCFormer’s low computational cost and fast inference time make it well suited for clinical applications requiring real-time responsiveness—such as automated triage, abnormality flagging, or preliminary report generation.
+With only 5.85 million parameters, the DCFormer-naïve variant can run on standard hospital hardware without requiring specialized infrastructure, significantly lowering the barrier to deploying advanced AI in routine workflows.
+We envision DCFormer as a foundational component in next-generation clinical decision support systems, helping democratize access to multimodal AI and accelerating the adoption of vision–language tools in healthcare.
+Beyond the CLIP framework, DCFormer has been independently validated as a versatile vision backbone in the recently proposed Med3DVLM architecture [33].
+In that work, DCFormer was combined with SigLIP [34] and incorporated into the LLaVA [35] framework to enable instruction-tuned multimodal understanding over 3D medical volumes and free-text prompts.
+The resulting system demonstrated competitive or superior performance in 3D report generation, zero-shot retrieval, and visual question answering tasks.
+This broader validation highlights DCFormer’s generalizability across different vision-language learning paradigms and tasks, reinforcing its potential as a robust foundation for multimodal medical AI.
+While DCFormer demonstrates strong potential, several avenues for future work remain.
+First, more sophisticated prompt engineering techniques could further improve zero-shot performance by better aligning the visual and textual embeddings for each clinical query.
+Second, incorporating additional and more diverse training data – including other imaging modalities like MRI and PET – could improve the model’s generalization and validate its effectiveness across different 3D imaging contexts.
+Finally, integrating DCFormer with emerging multimodal frameworks (such as LLaVA [35] for vision–language question answering or LISA [36] for segmentation) could broaden its application to tasks like interactive diagnosis and automated 3D image interpretation.
 Methods
-Processing 3D CT images presents significant computational challenges due to their high resolution and volumetric complexity.
-3D convolution is computationally expensive due to its cubic scaling with kernel size, and the self-attention mechanism further amplifies the burden by scaling quadratically with input size.
-To address these challenges, we propose DCFormer, a hybrid architecture that decomposes 3D convolution into multiple 1D components, significantly reducing computational overhead.
-DCFormer adopts MetaFormer [33, 36] as its backbone and employs decomposed convolutions as the token mixer for feature extraction.
-Its hierarchical structure further enhances feature representation in 3D CT images.
+Processing 3D images presents significant computational challenges due to their high resolution and volumetric complexity.
+The computational complexity of standard 3D convolutions scales cubically with kernel size, and self-attention mechanisms further amplify the burden by scaling quadratically with image size.
+To address these challenges, we developed DCFormer, a hybrid architecture that decomposes 3D convolutions into multiple 1D components, significantly reducing computational overhead.
+DCFormer adopts the MetaFormer paradigm [32, 37] as its backbone and employs decomposed convolutions as the token mixer for feature extraction.
+Its hierarchical design (described below) further enhances feature representation in 3D images.
 Finally, the DCFormer-based image encoder is integrated into a CLIP framework for image-text alignment.
-As demonstrated in our results, this decomposition strategy substantially reduces parameter count and computational cost while maintaining high performance in both zero-shot and fine-tuning settings.
-Formulation of DCFormer block
-The core concept, MetaFormer [33, 36], is a general architecture abstracted from transformers [37].
-In the Metaformer architecture, the token mixer (e.g.
+As demonstrated in our results, this decomposition strategy substantially reduces parameter count and computational cost while maintaining high performance.
+Formulation of the DCFormer block
+Our design builds on the MetaFormer architecture [32, 37], which is a general vision transformer paradigm.
+In the Metaformer framework, the token mixer (e.g.
 self-attention, depthwise convolution) is not specified, while normalization, a channel mixer (e.g.
 MLP), and residual connections [38] are retained.
-The input $I$ is first passed into a patch embedding block, such as a convolution:
+The input image tensor $I$ is first passed through a patch embedding block, such as a convolution:
 $X=\text{PatchEmbed}(I)=Flatten(\text{Conv}^{(k,s,p)}_{C_{0}\rightarrow C}(I))$ (1)
@@
 $X^{\prime}$ $=X+\text{Norm}_{h}(X^{*}_{\text{h}})+\text{Norm}_{w}(X^{*}_{\text{w}})+\text{Norm}_{d}(X^{*}_{\text{d}})$ (7)
-An illustration of the DCFormer and its Pytorch-like implementation are shown in Figure 3 and Algorithm 1, respectively.
+An illustration of the DCFormer block and its relationship to existing architectures is shown in Figure 3, and PyTorch-style pseudocode for the Decomposed Convolution operation is provided in Algorithm 1, respectively.
 Figure 3: Block illustration of MetaNeXt, ConvNeXt and DCFormer.
@@
 Complexity Analysis
-Given an input tensor $X\in\mathbb{R}^{1\times C\times H\times W\times D}$ with batch size $B=1$ a 3D depthwise convolution [29] with a kernel size ($k_{h},k_{w},k_{d}$) where $k_{h}=k_{w}=k_{d}=k$, the total parameter count and FLOPs can be calculated as $Ck^{3}$ and $2CHWDk^{3}$, respectively (bias is omitted for simplicity).
-Note that both parameters and FLOPS increase cubically as $k$ increases.
-This introduces extensive computational burden, especially for large kernel sizes and when these convolutions are used repeatedly in an end-to-end network, where the effect of the cubic scaling compounds over multiple layers.
-By decomposing the 3D depthwise convolution into three 1D depthwise convolutions, the parameter count and FLOPS are significantly reduced to $3Ck$ and $6CHWDk$, respectively, as each 1D convolution operates independently along a single spatial dimension rather than all three simultaneously.
-A comparison of FLOPs and parameter counts for 3D depthwise and decomposed convolutions is shown in Figure 2.
-For the decomposed convolution, the computational cost scales linearly with respect to kernel size $k$, as opposed to the cubic scaling in the 3D depthwise convolution.
-When the kernel size is larger than 3, the complexity increase becomes much more significant.
-Large kernel sizes such as 7, 11 [31, 32] and even up to 51 [43, 44] have been shown to improve model performance.
-In 3D imaging, our decomposition approach becomes crucial, particularly in deep networks where multiple convolution layers are applied.
-Figure 4: Overview of DCFormer
+Given an input tensor $X\in\mathbb{R}^{C\times H\times W\times D}$, a standard 3D depthwise convolution [29] with a kernel size $k$ in all dimensions has $Ck^{3}$ parameters and $2CHWDk^{3}$ FLOPs (bias operations are omitted for simplicity).
+This cubic growth in complexity can quickly become prohibitive for large $k$ and for deep networks, where multiple such layers compound the cost.
+In contrast, our decomposed 3D convolution uses three 1D depthwise convolutions, which collectively require only $3Ck$ parameters and $6CHWDk$ FLOPs.
+Thus, the parameter count and computational cost scale linearly with kernel size rather than cubically.
+Figure 2 compares the theoretical FLOPs and parameter counts of a standard 3D depthwise convolution with those of a decomposed convolution, highlighting the drastic efficiency gain of the latter.
+Notably, for any kernel size larger than 3, the difference in complexity becomes increasingly significant.
+Large kernel sizes (e.g., 7 or 11, as used in prior works [31, 43], and even 51 as explored in recent studies [44, 45]) have been shown to improve CNN performance.
+In 3D imaging, our decomposition approach makes the use of such large kernels feasible, which is particularly important in deep networks where multiple convolutional layers are stacked.
+Figure 4: Hierarchical architecture of DCFormer.
 Building a Vision Encoder with DCFormer
-Based on the above findings, we formulate DCFormer as the token mixer and develop a family of architectures of varying sizes, using DCFormer as the main building block.
-Figure 4 illustrates the general architecture of our proposed DCFormer.
-Following previous work [38, 45, 31, 33, 46, 32], DCFormer adopts a hierarchical structure with four stages and an initial stem stage following [46, 47].
+Using the DCFormer block as the core component, we constructed a hierarchical vision encoder architecture (Figure 4).
+Following prior work [38, 46, 31, 32, 47, 43], DCFormer includes an initial stem stage followed by four hierarchical stages.
 In the stem stage, we use a decomposed convolution with a kernel size of 7 and a stride of 4.
 The downsampled images are then processed with three additional decomposed convolutions, each with a kernel size of 3 and a stride of 1.
-The outputs of the stem stage are passed into the hierarchical structure, where each stage produces tokens of sizes $\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$, $\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$, $\frac{H}{32}\times\frac{W}{32}\times\frac{D}{32}$, and $\frac{H}{64}\times\frac{W}{64}\times\frac{D}{64}$, respectively, where $H$, $W$, $D$ represent height, width and depth of the input image volume.
-We propose four model variants: nano, naive, tiny, small.
-The number of layers in each stage are [1, 1, 1, 1] for nano, [2, 2, 2, 2] for naive, [2, 3, 3, 2] for tiny and [2, 3, 6, 2] for small models.
-Additionally, each stage begins with patch embedding where a 3D max pooling operation with a kernel size of 3 and a stride of 2 is used to reduce the image dimensions [47].
-Following [7, 31, 32], we set the MLP ratio to 4, while the kernel size for the decomposed convolution is set to 7, as in [45, 31].
-Further configuration details are provided in Table 2.
+The output of the stem is then fed into four stages of DCFormer blocks, where each stage produces tokens of sizes $\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$, $\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$, $\frac{H}{32}\times\frac{W}{32}\times\frac{D}{32}$, and $\frac{H}{64}\times\frac{W}{64}\times\frac{D}{64}$, respectively, where $H$, $W$, $D$ represent height, width and depth of the original image volume.
+We implemented DCFormer in three model variants: nano, naive, and tiny.
+The number of layers in each stage is [1, 1, 1, 1] for nano, [2, 2, 2, 2] for naive, and [2, 3, 3, 2] for tiny.
+At the beginning of each stage (except the stem), we include a patch embedding downsampling operation implemented via 3D max pooling with a kernel size of 3 and a stride of 2, to further reduce the feature map size [48].
+We set the MLP expansion ratio to 4 in all DCFormer blocks, following [7, 31, 43], and use a base decomposed convolution kernel size of 7, consistent with the design in [46, 31].
+Additional configuration details for each model variant are provided in Table 4.
+Table 4: Architectural configurations of the DCFormer model variants.
 Stage #Tokens Layer Specification DCFormer
-Nano Naive Tiny Small
+Nano Naive Tiny
 0 H/4 X W/4 X D/4 STEM Embed. Dim. 32 64
@@
 MLP Ratio 4
-#Blocks 1 2 3 6
+#Blocks 1 2 3
 4 H/64 X W/64 X D/64 Patch Embedding Patch size 3 x 3 x 3, stride 2
@@
 #Blocks 1 2
-Parameters (M) 0.92 5.85 15.1 26.86
-FLOPS(G) 34.21 49.48 168.2 202.9
-Table 2: DCFormer configuration details.
+Parameters (M) 0.92 5.85 15.1
+FLOPS(G) 34.21 49.48 168.2
 Developing a CLIP Framework using DCFormer
@@
 As a result, the DCFormer-based CLIP framework is well-suited for large-scale 3D medical imaging applications.
-Figure 5: CLIP a) training and b) zero-shot inference framework
-Zero-shot multi-abnormality detection with DCFormer
-For zero-shot multi-abnormality detection, we follow the same approach as CheXzero[21] and CT-CLIP[26].
-Specifically, two similarity scores are computed using the cosine similarity between the embedding of the CT image of interest and the embeddings of a positive (’{Pathology} is present.’) prompt and a negative (’{Pathology} is not present.’) prompt.
-A softmax function is then applied to these similarity scores to estimate the likelihood of each abnormality being present in the given CT image (see Figure.
-5 (b) for illustration).
-It should be noted that the zero-shot performance can potentially be improved by experimenting with different prompt variations.
-For instance, possible prompt templates including: ’Signs of {Pathology} are/are not detected in the image.’, ’There is/is not {Pathology}.’, ’{Pathology} is/is not observed.’ ’Indication/No indication of {Pathology}.’.
-Although prompt engineering plays a critical role in improving image-text alignment, we adopt the prompt structure proposed by Hamamci et al.
-[26] (’{ Pathology} is/is not present.’) to maintain consistency with prior work and ensure a fair comparison.
+Figure 5: DCFormer-based CLIP framework: (a) Training with paired CT volumes and reports, (b) Zero-shot inference with text prompts, (c) Fine-tuning for multi-label classification, and (d) Text-to-image retrieval based on embedding similarity.
 State-of-the-Art Image Encoders
@@
 Implementation Details
-For model training, we follow the same data splitting and pre-processing strategy as in [26].
-Specifically, we use 20,000 patients for training and 1,304 for validation.
-For pre-processing, we first resize each CT volume to a spacing of 0.75 mm on the x-axis and y-axis and 1.5 mm on the z-axis.
-Then, the volumes are center-cropped or padded to a fixed size of 512 x 512 x 256.
-Finally, we clip the Hounsfield Unit (HU) values of each CT image volume to [-1000,1000] and normalize them to [-1,1].
-We train models using the AdamW optimizer [48] with an initial learning rate $1^{-5}$.
-We dot not apply learning rate scheduling or warmup, as we have not observed any significant improvements.
-All models are trained for 15 epochs.
+For model training, we followed the same data splitting and pre-processing strategy as in [26].
+Specifically, we used 20,000 patients for training and 1,304 for validation.
+For pre-processing, we first resized each CT volume to a spacing of 0.75 mm on the x- and y-axes, and 1.5 mm on the z-axis.
+The volumes were then center-cropped or padded to a fixed size of 512 × 512 × 256.
+Finally, we clipped the Hounsfield Unit (HU) values of each CT image volume to the range $[-1000,1000]$ and normalized them to $[-1,1]$.
+We trained our models using eight NVIDIA A100 80GB GPUs.
+For optimization, we used AdamW [49] with an initial learning rate of $10^{-5}$.
+We did not apply learning rate scheduling or warmup, as we did not observe any significant improvement.
+All models were trained for 15 epochs.
+For fine-tuning, we integrated a linear classification layer into the image encoder and trained only this layer while keeping the image encoder frozen.
+Table 5: Detailed overview of the CT-RATE dataset for each abnormality and their distributions in the training and validation sets. [26]
 Abnormality Number of samples (train) Number of samples (validation) Ratio in the dataset (train) Ratio in the dataset (validation)
@@
 Interlobular septal thickening 3745 249 0.079 0.082
-Table 3: Detailed overview of the CT-RATE dataset for each abnormality and their distributions in the training and validation sets. [26]
 Acknowledgements
 This work was supported by the Department of Medicine and the Intelligent Clinical Care Center at the University of Florida College of Medicine.
-Author contributions statement
+The authors express their sincere gratitude to the NVIDIA AI Technology Center at the University of Florida for their invaluable feedback, technical guidance, and support throughout this project.
+Author Contributions
 G.C.A.
-conducted study design, software development, experimental setup, data analysis, visualization, and manuscript writing.
+conducted the study design, software development, experimental setup, data analysis, visualization, and manuscript writing.
+Y.X.
+contributed to the experimental setup, manuscript writing, and discussions.
 K.G.
-provided numerous insights and contributed to discussions throughout this study.
+provided valuable insights and contributed to discussions throughout the study.
 W.S.
-oversaw the project, conducted study design, experimental setup, and data analysis, and contributed to manuscript writing.
-All authors reviewed the manuscript.
+supervised and conceptualized the project, guided methodological development and experimental setup, and contributed to writing the original manuscript.
+All authors reviewed and approved the final manuscript.
 Additional Information
-Competing interests The authors have no conflict of interest to declare.
+Competing interests: The authors have no conflict of interest to declare.